# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### v2

@@ 5.1 Two Moons
The Two Moons model is a standard benchmark example in SBI (Greenberg et al., 2019; Lueckmann et al., 2021; Picchini & Tamborrino, 2022) to illustrate how algorithms deal with multimodality and the local crescent shapes of the posterior distribution. The model itself is defined in Appendix E. Our experiments are run on the same 10 datasets provided in SBIBM, each of them being a vector $\bm{y}_{0}$ of length two. We use $K=30$, a value we determine as described in Appendix B (this determination only required 90 seconds). We run SeMPLE, SNL and SNPE-C algorithms for a total budget of $10^{4}$ model simulations, distributed across inference rouces, and we tested both $R=4$ and $R=10$ rounds. The obtained samples are then compared with the corresponding posterior reference samples (obtained via MCMC, since the likelihood is tractable) provided in SBIBM. The performance metric results are averaged over the 10 different data sets. Figure 2 shows C2ST and runtimes values. SeMPLE provides the best approximation while SNL performs the worst. The comparison with SNL is particularly interesting as while both SNL and SeMPLE incorporate a tuning-free MCMC step, SeMPLE is able to visit both modes with an acceptance rate of 60-70%, while SNL often struggles in efficiently exploring a multimodal posterior as demonstrated by the higher C2ST values. Moreover, in this simple example the runtime for SeMPLE is about half the SNPE-C runtime, and is about three times lower than the SNL runtime. In other examples, (e.g. Figure 3), the speedup provided by SeMPLE is even more dramatic. All these benefits from SeMPLE come with a very small impact on the memory usage, see Section 5.4 for details.
@@ 5.2 Multiple hyperboloid model
Figure 3: Hyperboloid. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
@@ 5.3 Lotka-Volterra
The Lotka-Volterra predator prey model is a classical example of a population dynamics model. We consider the stochastic Markov jump process version of the model, which is defined in Appendix I. This model, although being relatively simple and characterized by a set of three interactions on the two species, each with a corresponding rate $\theta_{j}$ ($j=1,2,3$), encompasses many of the difficulties associated with larger, more complex systems. The observations consider the addition of non-negligible measurement error to the simulated dynamics, and the standard deviation $\sigma$ of this noise term is inferred as an additional parameter, resulting in a total of four model parameters to infer $(\theta_{1},\theta_{2},\theta_{3},\sigma)$. For this model, we are not primarily focusing on the performance comparison to other neural-based methods, as this version of the model, as a Markov jump process, is not included in SBIBM. Moreover, running times for a single model call are random even for repeated calls using the same parameter value. We use the Gillespie algorithm, as explained in Appendix I, and the number of “reactions” at any given $\bm{\theta}$ is stochastic. Due to the reasons above, the primary focus is to show that high quality inference can be obtained with SeMPLE using a comparably small computational effort, and compare our results against two versions of sequential Monte Carlo ABC (SMC-ABC). A “reference posterior” is obtained with an efficient version of SMC-ABC called blockedopt in Picchini & Tamborrino (2022), which is run for very many model simulations, around 3.3 million, see Appendix I for details. We also compare it with a more standard off-the-shelf SMC-ABC named standard. SeMPLE inference is obtained with only 30,000 model simulations. In Figure 5 we also include the corresponding inference obtained with around 30,000 model simulations for both blockedopt and standard SMC-ABC. With only 30,000 model simulations, SeMPLE manages to produce inference for all parameters similar to the inference returned after 3.3 million model simulations from blockedopt SMC-ABC, a 100-fold computational saving. We observe that, unlike SMC-ABC, SeMPLE produces accurate posterior inference for all model parameters, including the noise parameter $\sigma$, but with a much smaller number of model simulations compared to SMC-ABC. The latter fact, namely the SMC-ABC inefficiency in terms of model simulations, was not unexpected and known from Greenberg et al. (2019); Lueckmann et al. (2021). However it is reassuring to notice the excellent inference quality brought by SeMPLE in this challenging case study.
@@ 5.4 Resource requirement metrics
The metrics reported in Table 1 correspond to four models and have been measured in kiloJoules (kJ) using the PowerAPI package (Bourdon et al., 2013). The experiments were run on a machine with a CPU Intel® CoreTM i7-4790 CPU @ 3.60GHz and 16GB of DRAM. Memory cost and peak memory usage results have been measured by the tracemalloc Python library. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance. For instance, in the Two Moons case, the memory requirement (Peak Memory Usage) for SeMPLE is 13 times lower than SNPE-C and 27 times lower than SNL. The gain is even larger for the other models in Table 1. Similarly, for the Ornstein-Uhlenbeck model the energy consumption (CPU+DRAM) is more than 10 times smaller with SeMPLE.
@@ 5.5 Results summary
In Table 2 we give a summary of the main results from all experiments where an objective comparison with NN-based methods was possible, including those in the Appendices. Instead, for the Lotka-Volterra model we did not run NN-based methods and we reported comparisons with SMC-ABC both in section 5.3 and in Appendix I. Generally SeMPLE returns accurate inference, either by producing the smallest metric value (the smaller the better), as for Two Moons and Hyperboloid, or being second best after SNL, as for Bernoulli GLM and Ornstein-Uhlenbeck, also considering the larger variation of the SNPE-C results. However, even when SNL produces a smaller Wasserstein distance, it has to be balanced with respect to the running time. For Bernoulli GLM, SNL is between 18 and 75 times slower than SeMPLE, depending on the number of rounds, as SNL has to perform an expensive neural networks training at each round. For the same reason, when looking at Bernoulli GLM, SNPE-C is between 8 and 26 times slower than SeMPLE, again depending on how many times it needs to be retrained. When considering the Ornstein-Uhlenbeck model, the Wasserstein distances have a good degree of overlap between SNL and SeMPLE, but SNL is between 9.5 and 10.5 times slower than SeMPLE, while SNPE-C is between 17 and 22 times slower than SeMPLE. In Table 2 we highlight whether an experiment has a unimodal or a multimodal target, as this determines the type of metric we report. Namely we use the C2ST metric for multimodal targets and the Wasserstein distance for unimodal targets. The reasons for this are now discussed, however notice that in the Appendices we report results for both metrics in each experiment. In our experiments, we found the C2ST metric (advocated in Lueckmann et al., 2021) useful to evaluate the inference quality with multimodal targets, when draws from a reference posterior are available. However, we also found that at times C2ST produced contradicting results with unimodal targets. An example is the Binomial GLM model considered in Appendix G. There, it is obvious that the marginal posteriors returned by SeMPLE are accurate and that inference improved between the first two rounds, however C2ST showed that inference was apparently deteriorating for SeMPLE from round $r=1$ to $r=2$, which was not the case in reality. We also computed Wasserstein distances, which could confirm that the inference quality via SeMPLE did improve from round 1 to round 2. However, for multimodal targets, it is true that Wasserstein distances may not be a suitable tool, see e.g. Balaji et al. (2019). We therefore warn the reader from blindly relying on a single performance metric. It is also important to examine marginal and joint posterior densities.
@@ 6 Conclusion and perspectives
We have addressed the challenging question of Bayesian inference in the absence of explicit likelihoods. We have investigated mixture models through the GLLiM method, and shown that their expressivity could account for multivariate distributions, while their much simpler structure made them more amenable to efficient learning than NN solutions. The resulting new inference strategy named SeMPLE yielded inference that was on par with inference from the state-of-art neural-network (NN)-based methods that we examined (SNL and SNPE-C), while being more energy and memory efficient. For the considered examples, SeMPLE produced inference of superior or similar quality than NN-based methods, and when it did occasionally under-perform, it was by a negligible amount that can be largely be compensated by the much smaller runtime. In fact, in all tested cases, SeMPLE displayed much lower running times and memory requirements. The energy gains appear to increase with the complexity of the model simulator, and we expect to obtain even larger gains with real-life simulators. For example, for the Ornstein-Uhlenbeck model, SeMPLE achieved a 17-fold runtime speedup compared to SNPE-C. Although very preliminary, we hope this attempt, to meet the constraints of Green AI (Schwartz et al., 2020), will open the way to the development of more methods that can achieve accurate posterior inference, balancing the environmental impact of growing energy costs with the obtained results quality.
@@
Some open questions, that can be addressed in future research, follow. SeMPLE can benefit from the summarization of data when the dimension of $\bm{y}$ is large, just like other SBI methods. In addition to reducing the dimension, using summary statistics (as we did with the Lotka-Volterra model) can also improve the GLLiM performance. Indeed the performance of SeMPLE depends on the availability of good first approximations of the posterior and likelihood, which themselves depend on the ability of GLLiM to find probabilistic mappings between parameters and observations, in a regression-like manner using a Gaussian mixture structure. The construction of informative summary statistics is an important, but independent, research question that affects not only standard ABC methods, but also SBI methods based of neural networks, e.g. Chen et al. (2021); Radev et al. (2023b). Finally, following the discussion in Section 5.5, we wish to emphasize that measuring the quality of posterior inference in SBI is certainly an important question deserving further research, and that multiple metrics should be considered when evaluating experiments.
@@ Appendix B BIC computation for selecting $K$
In GLLiM, the number $K$ of components in the Gaussian mixture model has to be specified prior to performing the EM procedure. A too large value of $K$ may result in overfitting and unnecessary computational effort, while a too small value of $K$ may limit the ability to represent the relationship between $\bm{\theta}$ and $\bm{y}$. The Bayesian Information Criterion (BIC) is used in Deleforge et al. (2014) to guide the selection of $K$. It is given in equation (11)
@@
The GLLiM implementation in the xLLiM package removes a mixture component $k$ if the mixture probability becomes zero. Empirically it was found that when mixture probabilities becomes very small, GLLiM could occasionally crash with errors about failed internal matrix inversions. To prevent this, a threshold was set to remove mixture components with probabilities $\tilde{\eta}_{k}(\bm{\theta})$ below this threshold, which we set to be 0.005 for the multiple hyperboloid and Ornstein-Uhlenbeck models, zero for Two Moons (hence no component was deleted) or 0.03 for Bernoulli GLM and SLCP models. This check was repeated in every round of SeMPLE. This shows the need for setting reasonable initial number of mixture components $K$, as setting a too large value inflates the problem of small mixture component probabilities. However, in several cases our inference tasks worked out-of-the-box by setting the initial value to $K=20$ or 30.
@@ Appendix C SNPE-A
In SeMPLE we do not have such restrictions, meaning that (as in SNL) we can consider any prior $p(\bm{\theta})$, however this comes at the price of introducing MCMC sampling (as in SNL). However, unlike SNPE-A, in SeMPLE we are not constrained to reuse the estimated one-component Gaussian and replicate this $K$ times to obtain a $K$-components mixture, instead the GLLiM algorithm automatically determines the means and covariance matrices for each of the $K$ components, thus adding flexibility to the shapes of the posterior distributions that can be targeted.
@@ E.1 Two moons model definition
As in Greenberg et al. (2019) we set uniform priors $\theta_{1}\sim\mathcal{U}(-1,1)$, $\theta_{2}\sim\mathcal{U}(-1,1)$ By construction, this toy example has a posterior distribution with two crescent shapes for a given observation.

### v3

@@ 5.1 Two Moons
The Two Moons model is a standard benchmark example in SBI (Greenberg et al., 2019; Lueckmann et al., 2021; Picchini & Tamborrino, 2024) to illustrate how algorithms deal with multimodality and the local crescent shapes of the posterior distribution. The model itself is defined in Appendix E. Our experiments are run on the same 10 datasets provided in SBIBM, each of them being a vector $\bm{y}_{0}$ of length two. We use $K=30$, a value we determine as described in Appendix B (this determination only required 90 seconds). We run SeMPLE, SNL and SNPE-C algorithms for a total budget of $10^{4}$ model simulations, uniformly distributed across inference rounds, and we used both $R=4$ and $R=10$ rounds. The obtained samples are then compared with the corresponding posterior reference samples (obtained via MCMC, since the likelihood is tractable) provided in SBIBM. The performance metric results are averaged over the 10 different data sets. Figure 2 shows C2ST and runtime values. SeMPLE provides the best approximation while SNL performs the worst. The comparison with SNL is particularly interesting as while both SNL and SeMPLE incorporate a tuning-free MCMC step, SeMPLE is able to visit both modes with an acceptance rate of 60-70%, while SNL often struggles in efficiently exploring a multimodal posterior as demonstrated by the higher C2ST values. Moreover, in this simple example the runtime for SeMPLE is about half the SNPE-C runtime, and is about three times lower than the SNL runtime. In other examples, (e.g. Figure 3), the speedup provided by SeMPLE is even more dramatic. All these benefits from SeMPLE come with a very small impact on the memory usage, see Section 5.5 for details.
@@ 5.2 Multiple hyperboloid model
Figure 3: Hyperboloid. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with the same data vs the number of model simulations. Shaded bands enclose the min and max values.
@@ 5.3 Lotka-Volterra
The Lotka-Volterra predator prey model is a classical example of a population dynamics model. We consider the stochastic Markov jump process version of the model, which is defined in Appendix I. This model, although being relatively simple and characterized by a set of three interactions on the two species, each with a corresponding rate $\theta_{j}$ ($j=1,2,3$), encompasses many of the difficulties associated with larger, more complex systems. The observations consider the addition of non-negligible measurement error to the simulated dynamics, and the standard deviation $\sigma$ of this noise term is inferred as an additional parameter, resulting in a total of four model parameters to infer $(\theta_{1},\theta_{2},\theta_{3},\sigma)$. For this model, we are not primarily focusing on the performance comparison to other neural-based methods, as this version of the model, as a Markov jump process, is not included in SBIBM. Moreover, running times for a single model call are random even for repeated calls using the same parameter value. We use the Gillespie algorithm, as explained in Appendix I, and the number of “reactions” at any given $\bm{\theta}$ is stochastic. Due to the reasons above, the primary focus is to show that high quality inference can be obtained with SeMPLE using a comparably small computational effort, and compare our results against two versions of sequential Monte Carlo ABC (SMC-ABC). A “reference posterior” is obtained with an efficient version of SMC-ABC called blockedopt in Picchini & Tamborrino (2024), which is run for very many model simulations, around 3.3 million, see Appendix I for details. We also compare it with a more standard off-the-shelf SMC-ABC named standard. SeMPLE inference is obtained with only 30,000 model simulations. In Figure 5 we also include the corresponding inference obtained with around 30,000 model simulations for both blockedopt and standard SMC-ABC. With only 30,000 model simulations, SeMPLE manages to produce inference for all parameters that is similar to the inference returned after 3.3 million model simulations from blockedopt SMC-ABC, a 100-fold computational saving against the most efficient version of SMC-ABC. We observe that, unlike SMC-ABC, SeMPLE produces accurate posterior inference for all model parameters, including the noise parameter $\sigma$, but with a much smaller number of model simulations compared to SMC-ABC. The latter fact, namely the SMC-ABC inefficiency in terms of model simulations, was not unexpected and known from Greenberg et al. (2019); Lueckmann et al. (2021). However it is reassuring to notice the excellent inference quality brought by SeMPLE in this challenging case study.
@@
5.4 Biological model of the translation kinetics after mRNA transfection
We consider a two-dimensional stochastic differential equation (SDE) model of the translation kinetics after mRNA transfection, as studied in Pieschner et al. (2022). Only one of the two coordinates of the process is observed (with measurement noise). Details about the model are in Appendix K. Pieschner et al. (2022) perform several experiments, where data are either exact observations of the system or are perturbed with measurement error, and inferences for both an ODE and an SDE model are compared. However, here we only study the SDE model with data perturbed with measurement error. The parameter to infer is (we consider parameters on log-scale, same as for their priors) $\theta=(\log\delta,\log\gamma,\log k,\log m_{0},\log\textrm{scale},\log t_{0},\log\textrm{offset},\log\sigma)$. We ran inference for five datasets, each consisting of a time-series of length 60, each generated with different parameter values. Details about data-generating parameters, and other settings, are in Appendix K. For one of the five datasets, in Figure 6 we show the marginal posteriors from the third round of SeMPLE, in comparison to the parameter priors, while in Appendix K we also show comparisons with SNL and SNPE-C. For all datasets, SeMPLE accurately infers $(\delta,\gamma,t_{0},\textrm{offset},\sigma)$, and we confirm the findings in Pieschner et al. (2022) (which they obtain using exact Bayesian inference via the Hamiltonian Monte Carlo method implemented in Stan), namely that $k$, $m_{0}$ and scale suffer from identifiability issues (see the supplementary section A.4.1 in Pieschner et al., 2022 where their parameter $\theta_{2}$ is our $k$). In Appendix K we show that only SeMPLE can identify $t_{0}$ accurately, while SNL and SNPE-C provide a biased estimation of $t_{0}$. Moreover, SNPE-C is often unable to infer the measurement error variability $\sigma$, while this is not a problem for SeMPLE.
Figure 6: Translation kinetics model: in green are the marginal posteriors from SeMPLE (at round $r=3$), priors (solid black lines) and ground truth parameters (vertical red lines).
@@ 5.5 Resource requirement metrics
The metrics reported in Table 1 correspond to four models and have been measured in kiloJoules (kJ) using the PyJoules [^5] https://pyjoules.readthedocs.io/en/latest/ module found in the PowerAPI package (Bourdon et al., 2013) in Python, which makes use of the Running Average Power Limit (RAPL) technology available on Intel® CPUs. It is important to note that these energy measurements are based on hardware readings, and not on estimations depending on eg running time. The experiments were run on a machine with a CPU Intel® CoreTM i7-4790 CPU @ 3.60GHz and 16GB of DRAM. Memory cost and peak memory usage results have been measured by the tracemalloc Python library. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance. For instance, in the Two Moons case, the memory requirement (Peak Memory Usage) for SeMPLE is 13 times lower than SNPE-C and 27 times lower than SNL. The gain is even larger for the other models in Table 1. Similarly, for the Ornstein-Uhlenbeck model the energy consumption (CPU+DRAM) is more than 10 times smaller with SeMPLE.
@@ 5.6 Results summary
In Table 2 we give a summary of the main results from all experiments where an objective comparison with NN-based methods was possible, including those in the Appendices. Instead, for the Lotka-Volterra model we did not run NN-based methods and we reported comparisons with SMC-ABC both in section 5.3 and in Appendix I. Generally, SeMPLE returns accurate inference, either by producing the smallest metric value (the smaller the better), as for Two Moons and Hyperboloid, or being second best after SNL, as for Bernoulli GLM and Ornstein-Uhlenbeck, also considering the larger variation of the SNPE-C results. However, even when SNL produces a smaller Wasserstein distance, it has to be balanced with respect to the running time. For Bernoulli GLM, SNL is between 18 and 75 times slower than SeMPLE, depending on the number of rounds, as SNL has to perform an expensive neural network training at each round. For the same reason, when looking at Bernoulli GLM, SNPE-C is between 8 and 26 times slower than SeMPLE, again depending on how many times it needs to be retrained. When considering the Ornstein-Uhlenbeck model, the Wasserstein distances have a good degree of overlap between SNL and SeMPLE, but SNL is between 9.5 and 10.5 times slower than SeMPLE, while SNPE-C is between 17 and 22 times slower than SeMPLE. In Table 2 we highlight whether an experiment has a unimodal or a multimodal target, as this determines the type of metric we report. Namely we use the C2ST metric for multimodal targets and the Wasserstein distance for unimodal targets. The reasons for this are now discussed, however notice that in the Appendices we report results for both metrics in each experiment. In our experiments, we found the C2ST metric (advocated in Lueckmann et al., 2021) useful to evaluate the inference quality with multimodal targets, when draws from a reference posterior are available. However, we also found that at times C2ST produced contradicting results with unimodal targets. An example is the Bernoulli GLM model considered in Appendix G. There, it is obvious that the marginal posteriors returned by SeMPLE are accurate and that inference improved between the first two rounds, however C2ST showed that inference was apparently deteriorating for SeMPLE from round $r=1$ to $r=2$, which was not the case in reality. We also computed Wasserstein distances, which could confirm that the inference quality via SeMPLE did improve from round 1 to round 2. However, for multimodal targets, it is true that Wasserstein distances may not be a suitable tool, see e.g. Balaji et al. (2019). We therefore warn the reader from blindly relying on a single performance metric. It is also important to examine marginal and joint posterior densities. For example, in Bischoff et al. (2024) it is mentioned that interpreting results for C2ST requires knowledge of the classifier used and its appropriateness for the data at hand, and that its effectiveness is critically dependent on the selection and training of a suitable classifier.
@@ 6 Conclusion and perspectives
We have addressed the challenging question of Bayesian inference in the absence of explicit likelihoods. We have investigated mixture models through the GLLiM method, and shown that their expressivity could account for multivariate distributions, while their much simpler structure made them more amenable to efficient learning than NN solutions. The resulting new inference strategy, named SeMPLE, yielded inference that was on par with inference from the state-of-art neural-network (NN)-based methods that we examined (SNL and SNPE-C), while being more energy and memory efficient. For the considered examples, and the considered simulation setup, SeMPLE produced inference of superior or similar quality than NN-based methods, and when it did occasionally under-perform, it was by a negligible amount that can largely be compensated by the much smaller runtime. In fact, in all tested cases, SeMPLE displayed much lower running times and memory requirements. The energy gains appear to increase with the complexity of the model simulator, and we expect to obtain even larger gains with real-life simulators. For example, for the Ornstein-Uhlenbeck model, SeMPLE achieved a 17-fold runtime speedup compared to SNPE-C. Although very preliminary, we hope with this attempt to meet the constraints of “Green AI” (Schwartz et al., 2020), and open the way to the development of more methods that can achieve accurate posterior inference, balancing the environmental impact of growing energy costs with the obtained results quality.
@@
Some open questions, that can be addressed in future research, follow. SeMPLE can benefit from the summarization of data when the dimension of $\bm{y}$ is large, just like other SBI methods. In addition to reducing the dimension, using summary statistics (as we did with the Lotka-Volterra model) can also improve the GLLiM performance. Indeed the performance of SeMPLE depends on the availability of good first approximations of the posterior and likelihood, which themselves depend on the ability of GLLiM to find probabilistic mappings between parameters and observations, in a regression-like manner using a Gaussian mixture structure. The construction of informative summary statistics is an important, but independent, research question that affects not only standard ABC methods, but also SBI methods based of neural networks, e.g. Chen et al. (2021); Radev et al. (2023b). Finally, following the discussion in Section 5.6, we wish to emphasize that measuring the quality of posterior inference in SBI is certainly an important question deserving further research, and that multiple metrics should be considered when evaluating experiments. For a recent study on metrics to evaluate the quality of SBI, see Bischoff et al. (2024) .
@@ Acknowledgements
  11. K
    Biological model of the translation kinetics after mRNA transfection K
    1. K.1
      Model definitionK.1
    2. K.2
      Results K.2
@@ Appendix B BIC computation for selecting $K$
In GLLiM, the number $K$ of components in the Gaussian mixture model has to be specified prior to performing the EM procedure. A too large value of $K$ may result in overfitting and unnecessary computational effort, while a too small value of $K$ may limit the ability to represent the relationship between $\bm{\theta}$ and $\bm{y}$. The Bayesian Information Criterion (BIC), used in Deleforge et al. (2014) to guide the selection of $K$, is given in equation (11)
@@
The GLLiM implementation in the xLLiM package removes a mixture component $k$ if the mixture probability becomes zero. Empirically it was found that when mixture probabilities becomes very small, GLLiM could occasionally crash with errors about failed internal matrix inversions. To prevent this, a threshold was set within SeMPLE to remove mixture components with probabilities $\tilde{\eta}_{k}(\bm{\theta})$ below this threshold, which we set to be 0.005 for the multiple hyperboloid, Ornstein-Uhlenbeck, Lotka-Volterra and the biological SDE model, zero for Two Moons (hence no component was deleted) or 0.03 for Bernoulli GLM and SLCP models. This check was repeated in every round of SeMPLE. This shows the need for setting reasonable initial number of mixture components $K$, as setting a too large value inflates the problem of small mixture component probabilities. However, in several cases our inference tasks worked out-of-the-box by setting the initial value to $K=20$ or 30.
@@ Appendix C SNPE-A
In SeMPLE we do not have such restrictions, meaning that we can consider any prior $p(\bm{\theta})$, however this comes at the price of introducing MCMC sampling (as in SNL). However, unlike for SNPE-A, in SeMPLE we are not constrained to reuse the estimated one-component Gaussian and replicate this $K$ times to obtain a $K$-components mixture, instead the GLLiM algorithm automatically determines the means and covariance matrices for each of the $K$ components, which as such can be different between components, thus adding flexibility to the shapes of the posterior distributions that can be targeted.
@@ E.1 Two moons model definition
Same as in Greenberg et al. (2019), we set uniform priors $\theta_{1}\sim\mathcal{U}(-1,1)$, $\theta_{2}\sim\mathcal{U}(-1,1)$ By construction, this toy example has a posterior distribution with two crescent shapes for a given observation.